# Mitigating photon loss in linear optical quantum circuits: classical postprocessing methods outperforming postselection

### v2

@@ Abstract
Photon loss rates set an effective upper limit on the size of computations that can be run on current linear optical quantum devices. We present a family of techniques to mitigate the effects of photon loss on both output probabilities and expectation values derived from noisy linear optical circuits composed of an input of $n$ photons, an $m$–mode interferometer, and $m$ single photon detectors. Central to these techniques is the construction of objects called recycled probabilities. Recycled probabilities are constructed from output statistics affected by loss, and are designed to amplify the signal of the ideal (lossless) probabilities. Classical postprocessing techniques then take recycled probabilities as input and output a set of loss-mitigated probabilities, or expectation values. We provide analytical and numerical evidence that these methods can be applied, up to large sample sizes, to produce more accurate outputs than those obtained from postselection - which is currently the standard method of coping with photon loss when sampling from discrete variable linear optical quantum circuits. In contrast, we provide strong evidence that the popular zero noise extrapolation technique cannot improve on on the performance of postselection for any photon loss rate.
@@ 1 Introduction
Discrete variable linear optical quantum computing (DVLOQC) is a framework that uses a discrete number of photons as well as linear optical hardware to store and process quantum information. Many models of universal and fault-tolerant quantum computation tailored to this framework have been developed over the years beginning with the work of [1], and followed by various other models and variants (e.g [2, 3, 4, 5]). Furthermore, promising proposals for the near-term demonstration of quantum-over-classical advantage such as boson sampling [6] can naturally be implemented in this framework. A quantum device capable of performing DVLOQC usually consists of three components: single photon sources [7], multimode interferometers [8], and single photon detectors [9]. We will refer to the collection of these as a linear optical circuit. One major impediment to scaling up such a device is photon loss [10]. Postselection, where all lossy output statistics with one or more lost photons are discarded and only those where all photons are detected are kept, may be used to obtain the ideal output distribution of a linear optical circuit subject to photon loss [11]. This can be viewed as a form of quantum error mitigation, where the ideal distribution is accessible, but with a sampling cost scaling exponentially with the depth of the circuit. This scaling comes from the fact that the probability of at least one photon being lost approaches one exponentially quickly with increasing circuit depth [12], which results in most outputs being discarded.
@@
In this work we address the question of whether the effects of photon loss in the output distributions of linear optical quantum circuits can be mitigated by classical postprocessing of lossy output statistics. We provide a positive answer to this question by presenting various new techniques to mitigate photon loss in linear optical circuits, and performing a rigorous analysis of their performance. We refer to these techniques collectively as recycling mitigation, as they all involve the use of lossy output statistics that otherwise would be discarded. At the heart of recycling mitigation is the construction of the so-called recycled probabilities. The recycled probabilities can be thought of as precursors of the mitigated outputs. Mitigated probabilities are approximations of the ideal probabilities, and can be obtained from the recycled probabilities by appropriate classical postprocessing. We introduce several techniques by which this classical postprocessing step can be performed. This in turn gives rise to different photon loss mitigation techniques. Fig. 1 shows the main steps underlying our mitigation techniques.
We provide strong analytical and numerical evidence that in the high loss regime where the probability, denoted as $\eta$, that a photon is lost in any given mode satisfies $\eta>0.5$, recycling mitigation outperforms postselection. The intuition being that in this regime, output events where photons are lost are much more likely to occur than events where no photon is lost. Since recycling mitigation uses lossy output statistics, it is natural to expect that the resulting mitigated probabilities are more converged than probabilities computed only from the postselected samples. Furthermore, we provide analytic and numerical evidence showing that mitigation techniques based on artificially increasing noise and Richardson extrapolation, called zero noise extrapolation (ZNE) [13, 14], such as those applied to mitigate photon loss in the continuous variable regime [15], offer no advantage over postselection for the problem of mitigating photon loss in the DVLOQC setting.
Recycling mitigation can be applied to linear optical circuits implementing arbitrary unitaries. In particular, when these unitaries are randomly chosen according to the Haar measure, this gives rise to the non-universal model of computation known as boson sampling [6]. There are a number of classical algorithms for boson sampling that become efficient when photon loss is sufficiently high [16, 12]. In this spirit, we show that recycled probabilities constructed from lossy output statistics where $k$ out of $n$ photons are lost, and $n-k$ is a constant independent of $n$, can be efficiently computed classically. These correspond to lossy output statistics where the majority of photons are lost. Conversely, and using results from [17], we provide evidence that recycled probabilities constructed from output statistics where $k$ is a constant independent of $n$ are hard to compute classically in worst-case. This corresponds to the case of output statistics where few photons are lost. These results indicate that recycling mitigation provides interesting performances when looking at output statistics with a small number of lost photons, although there is a set of intermediate lossy output statistics, where $k$ does not follow either of the above conditions, that we do not explore, and that might also give interesting performances.
Crucially, our developed methods are applicable to current and near-term photonic quantum hardware, and, as will be seen later, provide better performances in the regime of high loss ($\eta>0.5$). This is in contrast with fault-tolerant photonic quantum computing [3, 4] which typically requires large numbers of photons, as well as very low loss levels ($\eta\ll 0.5$), both of which are beyond the reach of currently available hardware. Furthermore, as discussed in the previous paragraph, the output statistics used in constructing the mitigated probabilities should be those where the number of lost photons $k$ is low. Thus, these conditions define the realm of applicability of recycling mitigation that provides the best performance: Recycling mitigation should be applied in the $\eta>0.5$ regime, and the output statistics used in constructing the mitigated probabilities should be those where $k$ is low.
Recycling mitigation can be used in a variety of applications, namely any DVLOQC algorithm where the desired output is an expectation value of an observable, a set output of probabilities, or a full probability distribution. Applications include boson sampling [6], photonic variational quantum eigensolvers [18, 19], photonic differential equation solving [20], photonic quantum machine learning [21], and graph problems with DVLOQC [22]. In a related article [23], we develop a quantum circuit born machine (QCBM) [24] tailored to DVLOQC. We also demonstrate significant photon loss mitigation, using recycling mitigation, for the training of the QCBM in various scenarios.
@@
This paper is structured as follows. Section 2 gives a high-level overview of our main contributions. Section 3 compares our techniques with those existing in the literature. Section 4 introduces some notation and basic concepts. Sections 5-8 detail the construction of the recycled distributions, the classical postprocessing techniques needed to obtain the mitigated distribution, as well as derive analytical conditions for recycling mitigation techniques to outperform postselection. Sections 9 and 10 contain numerical simulations that aid in understanding how to use our developed mitigation techniques in practice, as well as examples of our techniques in action. Section 11 provides strong evidence that techniques based on ZNE do not in general outperform postselection. Section 12 discusses how our techniques can be used to mitigate both expectation values of observables as well as full distributions. Section 13 discusses ways to improve our performance guarantees, and contains a technical result about sums of permanents of i.i.d. Gaussian matrices [6] that might find use beyond this work. Finally, section 14 presents a set of interesting open questions.
@@ 2 Overview of main results
In this section, we give a high-level overview of our key results, which are expanded upon in later parts of the paper.
Our first contribution is the construction of the recycled probabilities, which are the input of all our developed loss mitigation techniques. Consider an ideal $n$–photon probability $p_{id}(s)$, corresponding to a detection pattern $s$ of $n$ input photons passed through an $m$-mode linear optical interferometer, and detected in $m$ output modes. Its corresponding recycled probability $p^{k}_{R}(s)$ is an $n$–photon probability computed from output statistics affected by loss. Namely, $p^{k}_{R}(s)$ is constructed as a sum of of $n-k$–photon probabilities where $k$ out of the initial $n$ photons are lost. The precise construction and choice of the $n-k$–photon probabilities for each $s$ are detailed in section 5. Using the assumption of uniform loss ($\eta$ is the same in all modes) to express the $n-k$–photon probabilities as a convex sum of $n$ photon probabilities, we show that for all $s$ the recycled probabilities may be written in the form
$p^{k}_{R}(s)=\frac{1}{{m-n+k\choose k}}p_{id}(s)+\bigg{(}1-\frac{1}{{m-n+k\choose k}}\bigg{)}I_{s,k},$
where $I_{s,k}$ is a convex combination of unwanted $n$–photon probabilities.
In section 6, we show that recycled probabilities constructed from $n-k$–photon probabilities where $n-k$ is a constant independent of $n$ are efficiently computable classically. This motivates using $n-k$ output statistics where the value of $k$ is low relative to $n$ to construct recycled probabilities for use in recycling mitigation. One intuition for this is that higher-order photon statistics contain more information on the ideal $n$–photon distribution than the lower-order statistics. And there are results indicating that the ideal $n$–photon output distribution is hard to classically simulate, and further that simulating the $n-k$–photon output distribution is hard when $k$ is a constant [25, 17, 6].
The key idea behind all the techniques we develop for recycling mitigation is simple: try to extract $p_{id}(s)$ as accurately as possible from $p^{k}_{R}(s)$ via classical postprocessing. Since $p^{k}_{R}(s)$ can be obtained from the $n-k$ output photon statistics, if we had access to the value of $I_{s,k}$ we could directly compute $p_{id}(s)$ using the previous equation. Unfortunately it is not straightforward to compute the value of $I_{s,k}$, so instead we resort to deriving statistical inequalities bounding the variation of $I_{s,k}$ around its expected value. We derive these inequalities in the no-collision regime [6]. The details of this are found in section 7. Indeed, our second set of contributions begins with showing that the expected value $\mathbf{E}_{s}(I_{s,k})$ of $I_{s,k}$ over all $s$ is
$\mathbf{E}_{s}(I_{s,k})=\frac{1}{{m\choose n}}.$
This holds for arbitrary $m$-mode linear optical interferometers. Then, we use statistical inequalities [26, 27] to show that, for arbitrary $m$-mode interferometers, and all $s$,
$\epsilon_{bias,s}:=|I_{s,k}-\mathbf{E}_{s}(I_{s,k})|\leq\frac{1}{\mathsf{poly}(m)}.$
The above equation holds, for a given $s$, with probability $1-O\Big{(}\frac{\mathsf{poly}(m)}{{m\choose n}}\Big{)}.$ In addition to this probabilistic bound, for a broad family of $m$-mode interferometers having a specific structure, and using a result of [28], we manage to deterministically bound $\epsilon_{bias,s}$ as
$\epsilon_{bias,s}\leq O(e^{-0.000002n}).$
Bounding the bias error term $\epsilon_{bias,s}$ is essential for the later performance analyses of the techniques.
Our third set of contributions regards the development of classical postprocessing techniques for extracting estimates of $\{p_{id}(s)\}$ from $\{p^{k}_{R}(s)\}$, which we call the mitigated values $\{p_{mit}(s)\}$. These classical postprocessing techniques can be used to mitigate a full probability distribution of $n$ photons, a subset of the full distribution, or an expectation value of an observable $\mathsf{O}$, $\langle\mathsf{O}\rangle:=\sum_{i}w_{i}p_{id}(s_{i})$, with $w_{i}\in\mathbb{R}$, which is a weighted sum over a set of $n$–photon probabilities. The main cost incurred in these classical postprocessing techniques is a memory cost proportional to the size of the set of probabilities to be mitigated. In the worst-case, when we are mitigating a full probability distribution, this memory cost is $O\big{(}{m\choose n}\big{)}$. In practice, however, when mitigating an expectation value one usually needs access to a $\mathsf{poly}(n)$ (often of low degree) sized set of mitigated probabilities, making the memory cost manageable, and consequently our methods practical to apply to near-term DVLOQC hardware. Details about these classical postprocessing techniques can be found in section 8, and pseudocode for their implementation is given in algorithms 2, 3, 4, and 5. Note that since recycling mitigation is applied to the statistics relating to relatively few lost photons, the number of samples (runs of the lossy DVLOQC circuit) required for constructing the recycled probabilities, and computing the mitigated values to a fixed accuracy, scales exponentially with system size. On the other hand, this is a limitation common to all error mitigation protocols involving solely classical postprocessing [29, 30, 31], and not just specific to our technique.
In section 8 we provide performance guarantees for the classical postprocessing techniques. More precisely, for several of the postprocessing methods we bound the deviation of $p_{mit}(s)$ from $p_{id}(s)$ in terms of the bias error $\epsilon_{bias,s}$, as well as the statistical error $\epsilon_{stat,k}$ incurred when constructing the recycled probabilities from a finite number of samples. We observe that, in the $\eta>0.5$ regime, and up to a certain number of samples $N_{max}$, the combined worst-case bias and statistical errors of recycling mitigation are lower than the statistical error of postselection. This is strong evidence that our error mitigation methods outperform postselection in this regime. Furthermore, we observe through various numerical simulations in sections 9-10, that $N_{max}$ is large, meaning our methods are expected to outperform postselection for up to large sample sizes.

### v3

@@ Abstract
Photon loss rates set an effective upper limit on the size of computations that can be run on current linear optical quantum devices. We present a family of techniques designed to mitigate the effects of photon loss on both output probabilities and expectation values derived from noisy linear optical circuits composed of an input of $n$ photons, an mm–mode interferometer, and $m$ single photon detectors. Central to these techniques is the construction recycled probabilities. Recycled probabilities are constructed from output statistics affected by loss, and are designed to amplify the signal of the ideal (lossless) probabilities. Classical postprocessing techniques then take recycled probabilities as input and output a set of loss-mitigated probabilities, or expectation values. Our postprocessing methods result in biased estimators of the lossless probabilities. Nevertheless, we provide both analytical and numerical evidence that these methods can be applied, up to large sample sizes, to produce output probabilities with lower combined bias and statistical errors than the statistical errors of the output probabilities obtained from postselection. Therefore, these methods can outperform postselection - currently the standard method of coping with photon loss when sampling from discrete variable linear optical quantum circuits. In contrast, we provide evidence that the popular zero-noise extrapolation technique cannot improve on the performance of postselection for any photon loss rate.
@@ 1 Introduction
Discrete variable linear optical quantum computing (DVLOQC) is a framework that uses a discrete number $n$ of photons as well as $m$-mode linear optical interferometers to store and process quantum information. Many models of universal and fault-tolerant quantum computation tailored to this framework have been developed over the years beginning with the work of [31], and followed by various other models and variants (e.g [51, 14, 6, 52]). Furthermore, promising proposals for the near-term demonstration of quantum-over-classical advantage such as boson sampling [1] can naturally be implemented in this framework. A quantum device capable of performing DVLOQC usually consists of three components: single photon sources [58], multimode interferometers [53], and single photon detectors [22]. We will refer to the collection of these as a linear optical circuit. One major impediment to scaling up such a device is photon loss [70]. Postselection, where all lossy output statistics with one or more lost photons are discarded and only those where all photons are detected are kept, may be used to obtain the ideal output distribution of a linear optical circuit subject to photon loss [12]. This can be viewed as a form of quantum error mitigation, where the ideal distribution is accessible, but with a sampling cost scaling exponentially with the depth of the circuit. This scaling comes from the fact that the probability of at least one photon being lost approaches one exponentially quickly with increasing circuit depth [54], which results in most outputs being discarded.
@@
In this work we address the question of whether the effects of photon loss in the output distributions of linear optical quantum circuits can be mitigated by classical postprocessing of lossy output statistics. Crucially, we require that these postprocessing techniques outperform postselection, in the sense that they provide more converged estimates of the ideal probabilities for a comparable sample number. We present various new techniques to mitigate photon loss in linear optical circuits, and perform rigorous analysis of their performance. We refer to these techniques collectively as recycling mitigation, as they all involve the use of lossy output statistics that otherwise would be discarded. At the heart of recycling mitigation is the construction of the so-called recycled probabilities. The recycled probabilities can be thought of as precursors of the mitigated outputs. Mitigated probabilities are approximations of the ideal probabilities, and can be obtained from the recycled probabilities by appropriate classical postprocessing. We introduce several techniques by which this classical postprocessing step can be performed. These give rise to different photon loss mitigation techniques. Fig. 1 shows, at a very high level, the main steps underlying our mitigation techniques. We provide analytical and numerical evidence that there exists a threshold value, lower bounded by a constant, for the loss per mode $\eta$, denoted as $\eta_{th}$, such that when $\eta\geq\eta_{th}$, recycling mitigation outperforms postselection.
Furthermore, we provide analytic and numerical evidence showing that mitigation techniques based on artificially increasing noise and Richardson extrapolation, called zero-noise extrapolation (ZNE) [65, 36], such as those applied to mitigate photon loss in the continuous variable regime [62], do not outperform postselection for the problem of mitigating photon loss in the DVLOQC setting. This result is particularly interesting because we also show that, in the case of photon loss, ZNE produces unbiased estimators of the ideal probability. Indeed, we show that the application of ZNE to mitigate photon loss reduces to the problem of inverting a Vandermonde matrix [20] which, in the absence of statistical error, perfectly computes the ideal probabilities. In the presence of statistical error, however, we show that the resulting error on the inversion process is higher than the statistical error of postselection. The requirement of outperforming postselection, which is itself an unbiased estimator of the ideal probability, is what motivated us to search for biased loss mitigation techniques with smaller combined bias and statistical error than postselection.
Recycling mitigation is a biased error mitigation technique, meaning that in addition to statistical error there is also a bias error which, contrary to statistical error, does not decrease with increasing sample size. Nevertheless, we provide strong evidence that: $(1)$ For $\eta\geq\eta_{th}$, there is a sample size up to which the combined statistical and bias errors of recycling mitigation are less than the statistical error of postselection, meaning that recycling mitigation outperforms postselection up to this sample size. $(2)$ We present analytic and numerical evidence that the sample size after which postselection starts to outperform recycling mitigation seems very large, i.e. of the order of ${m\choose n}^{2}$, where $m\choose n$} is the size of the Fock space. $(3)$ We provide analytic and numerical evidence that, for fixed $m$ and $n$, and generic interferometers, the bias error in some of our introduced techniques seems small enough such that, in the limit of very low statistical error the mitigated outputs, like the ideal outputs, seem hard to approximate efficiently classically (see Section 6).
Any loss parameter $0<\eta<1$ results in a binomial distribution over $k$, the number of lost photons at the output. The key observation underpinning the utility of recycling is that when $\eta\geq\eta_{th}$, the statistical error denoted $\epsilon^{stat}_{k}$ is greatest for the postselection estimators where $k=0$, and furthermore that $\epsilon^{stat}_{k=0}\geq\epsilon^{stat}_{k=1}\geq\ldots\geq\epsilon^{stat}_{k=n_{d}}$, where $k\in\{1,\dots,n_{d}\}$ denotes the number of photons lost. $\epsilon^{stat}_{k}$ is the statistical error of recycled probabilities constructed from statistics with $k$ lost photons. In order to obtain the mitigated outputs these techniques use sampled outputs for which $k>0$ meaning these lossy estimators have lower statistical error than the postselection estimators, but where $k$ is still low enough such that the estimators contain information on the ideal outputs. The fact that statistical errors in recycling mitigation are lower than those of postselection, coupled with the low bias error in our techniques, is what allows the combined bias and statisical errors of the mitigated outputs to be lower than the statisical error of postselection, and thus for our methods to outperform postselection.
Since recycling mitigation is applied to the statistics relating to relatively few lost photons, we show that the number of samples (runs of the lossy DVLOQC circuit) required for constructing a recycled probability at a given $k$, and consequently computing the mitigated values to a fixed accuracy, scales exponentially with system size as $O\Big(\frac{1}{{n\choose k}(1-\eta)^{n-k}\eta^{k}}\Big)$. On the one hand, this is a significant improvement over the scaling of postselection when $\eta\geq\eta_{th}$, which is typically $O\big(\frac{1}{(1-\eta)^{n}}\big)$. On the other hand, exponential scaling is a bottleneck common to all error mitigation protocols involving solely classical postprocessing [15, 63, 50, 67], and not one specific to our technique. Finally, the fact that error mitigation techniques have exponential scaling does not necessarily hinder the achievement of a quantum computational advantage or utility for a fixed system size [74]. Indeed, an initial experimental attempt to achieve a quantum computational utility for a pre-fault tolerant quantum device coupled with error mitigation was presented in [30]. This experiment was later shown to be efficiently simulable classically, albeit using a highly non-trivial tensor network approach [66]. We are hopeful that other attempts in this direction will soon appear. Notably, in the DVLOQC framework where the methods we present can potentially aid in reaching a quantum computational utility in the pre-fault tolerant era.
The photon loss mitigation techniques we introduce can be used in a variety of applications. The work presented in [57] introduced a quantum circuit Born machine (QCBM) [7] tailored to the DVLOQC framework. Performance during the training of the QCBM, when afflicted by photon loss, considerably improved when the recycling mitigation method was applied. Use of the mitigation in the QCBM experiments reduced the overall number of samples needed to reach a desired precision, thereby decreasing the overall computational time. In addition to the application mentioned in [57], other applications include variational quantum eigensolvers [40, 34], photonic differential equation solving [24], photonic quantum machine learning [18], and graph problems with DVLOQC [43].
Recent work [64] has provided a method for mitigating photon loss in the continuous variable (CV) setting, adapting pre-existing methods of probabilistic error cancellation [65, 61, 39, 68] to the CV setting. Although the techniques of [64] are applicable to the DVLOQC setting, these are solely for expectation value mitigation (sometimes called weak mitigation) [50]. Our methods, by contrast, can perform both strong (full probability distribution mitigation) as well as weak mitigation, with a classical memory cost scaling with the size of the set of probabilities to be mitigated. Furthermore, we present analytical and extensive numerical evidence that our methods provably outperform postselection, whereas, to our knowledge, no comparison between the performance of the mitigation methods developed and postselection is presented in [64].
@@
This paper is structured as follows. Section 2 gives a high-level overview of our main contributions. Section 3 introduces some notation and basic concepts. Sections 4.1-4.5 detail the construction of the recycled distributions, the classical postprocessing techniques needed to obtain the mitigated distribution. Section 5 contains numerical simulations that aid in understanding how to use recycling mitigation in practice, as well as examples of the techniques in action. Section 6 discusses the properties of the mitigated probabilities in the limit of zero statistical error. Section 7 provides strong evidence that techniques based on ZNE do not in general outperform postselection. Section 8 contains a discussion of our results as well as a set of interesting open questions.
@@ 2 Overview of main results
(a)                       (b)
Figure 1: A schematic illustrating the main steps of the recycling mitigation protocol. (a) An input state of $n$ photons is introduced to a lossy $m$-mode linear optical interferometer implementing a unitary transformation. The classical measurement outcomes of each of the output modes is denoted by a classical $m$-bit string. (b) The classical data set generated by repeatedly sampling from the quantum circuit is then used as input for classical postprocessing. This classical postprocessing consists of three stages. First, the data set is used to generate lossy probability estimators. These estimators are then used to construct recycled probability estimators, which are in turn then used to generate mitigated values.
This section summarises the main technical contributions of this work.
First, we compute the overall error on the output probabilities when applying the recycling mitigation technique. That is, we compute the difference between the mitigated and ideal (lossless) output probabilities. This overall error can mainly be seen as a sum of statistical error, as well as the bias error of the technique. This result is stated precisely in the following theorem.
Theorem 1.
Consider $N_{tot}$ samples generated from a DVLOQC circuit with $n$ photons and $m$ modes. Let $U\in\mathsf{U}(m)$ be the unitary implemented by the circuit , $\eta\in[0,1]$ the uniform probability a photon is lost in any mode. Assume we are in the no-collision regime, $m\in\Omega(n^{2})$, where at most one photon occupies each mode. There exists a classical algorithm (linear solving recycling mitigation) that uses a subset of
$N_{rec}={n\choose k}(1-\eta)^{n-k}\eta^{k}N_{tot}\in O\Big(\sqrt{\frac{n}{k(n-k)}}\;\frac{n^{n}}{k^{k}(n-k)^{\,n-k}}\eta^{k}(1-\eta)^{n-k}N_{tot}\Big),$
samples where $k>0$ photons were lost, runs in time $N_{rec}+\mathsf{poly}(m,n,k)$, and that outputs with probability $1-\delta$ an approximation $p_{mit}(s)$ of the exact lossless circuit probability of bit string $s$, $p_{id}(s)$ such that
$|p_{mit}(s)-p_{id}(s)|\leq f(n,m,k,\eta,N_{tot}),$ (1)
where $f(n,m,k,\eta,N_{tot})=3\Bigg[\sqrt{\frac{2}{\delta}}\left(\frac{e(m-n+k)}{k}\right)^{k}(\frac{n}{m})^{n/2}+\sqrt{\ln(4/\delta)}\left(\frac{k}{n}\right)^{k/2}\sqrt{\frac{1}{(1-\eta)^{n-k}\eta^{k}N_{\mathrm{tot}}}}\Bigg]$.
This theorem is proved in Appendix 12.2. The proof makes use of Theorem 10 to bound the bias error of the mitigated output, and as the derived bound does not have an explicit dependence on $k$ it seems likely this could be considerably tightened. As such, Theorem 1 provides a loose upper bound on $f$. For example, if one looks at minimising $f$ by varying $k$, one finds $k_{min}=\mathsf{ceil}(\eta n)$ minimises $f$. It is tempting to say that this is an optimal value of $k$ which should be used in the protocol. However, this is a misleading conclusion directly due to the upper bound for $f$ being loose. Indeed, numerically we find that the error mitigated probabilities for values of $k$ near $k_{min}$ are far from the ideal probabilities. Intuitively, this is because the recycled probabilities used to extract the error mitigated probabilities can efficiently be estimated classically at $k=k_{min}$ by collecting samples from a classical algorithm simulating lossy boson sampling. Indeed, collecting samples from a lossy boson sampler near the average number of lost photons, $\mathsf{ceil}(\eta n)$, can be done efficiently classically for all constant $\eta$ [54]. Numerically, we find that values of $k\in O(1)$ produce the most accurate mitigation results (see Section 5). Note that these are rare sampling events for the classical sampler of [54]. We believe that any hope of proving the optimality of $k\in O(1)$ values requires a significant tightening of the upper bound on $f$, which in turn requires a more detailed understanding of the distribution of permanents of random matrices, which to the best of our knowledge remains incomplete [46].
Despite its limitations, Theorem 1 is still useful to show that a regime of advantage over postselection exists, as we will now show.
Postselection uses a subset $N_{post}=(1-\eta)^{n}N_{tot}$ of samples where $k=0$ (i.e. no loss occurred) and outputs an estimate $p_{post}(s)$ of $p_{id}(s)$ such that
$|p_{post}(s)-p_{id}(s)|\leq\gamma\sqrt{\frac{1}{(1-\eta)^{n}N_{tot}}},$ (2)
with probability $1-2e^{-\gamma^{2}}$ and runtime $N_{post}+1$. This upper bound on the additive error is computed via Hoeffding’s inequality [25].
The regime of advantage for the mitigation technique over postselection occurs where the combined statistical and bias errors of the mitigated output are lower than the statistical errors of the postselected output. To see the sampling condition where the recycling mitigation outperforms postselection indicated by the analytical bounds one can, in the worst-case where the additive errors of recycling mitigation and postselection equal their upper bounds in eqn. (1) and (2), solve the inequality: $|p_{mit}(s)-p_{id}(s)|\leq|p_{post}(s)-p_{id}(s)|$. This is stated in the following corollary.
Corollary 2.
We first assume that we are operating in the worst-case case error regime where $|p_{mit}(s)-p_{id}(s)|=f(m,n,k,\eta,N_{tot})$ (c.f. Theorem 1) and $|p_{post}-p_{id}|=\gamma\sqrt{\frac{1}{(1-\eta)^{n}N_{tot}}}$. Furthermore, we assume that the statistical error of the postselected outputs is higher than the statistical error of the (nk)(n-k)-photon lossy outputs. Recycling mitigation outperforms postselection when
$N_{\mathrm{tot}}\leq\frac{\delta(\Delta\eta)^{2}}{18}\Big(\frac{em}{n}\Big)^{n}\left(\frac{k}{e(m-n+k)}\right)^{2k},$ (3)
where
$\Delta\eta:=\gamma\sqrt{\frac{1}{(1-\eta)^{n}}}-3\sqrt{\ln(\!\frac{4}{\delta})}\,\left(\frac{k}{n}\right)^{k/2}\sqrt{\frac{1}{(1-\eta)^{n-k}\eta^{k}}},$ (4)
and, from a union bound, this upper bound on $N_{tot}$ applies with confidence $1-\delta^{\prime}$, where $\delta^{\prime}=\delta+2e^{-\gamma^{2}}$.
Note that the assumption in the above corollary that the statistical error from postselection is higher than the statistical error of the mitigated output may be concisely stated as $\Delta\eta>0$. This corollary means that, in worst-case, there exists a non-trivial sampling regime up to which recycling mitigation outperforms postselection. Furthermore, the condition $\Delta\eta>0$ indicates that there is a photon loss rate threshold that needs to be satisfied in order for recycling mitigation to outperform postselection. Indeed, $\Delta\eta>0$ implies the existence of a lower bound on the photon loss rate for which it is possible for mitigation to outperform postselection: $\eta>\eta_{th}(n,k)$ with
$\eta_{th}(n,k)\geq\frac{1}{1+\frac{en}{k}}.$
For sufficiently large $n$ and for a fixed $k$ independent of $n$, we have that $\left(\frac{e(m-n+k)}{k}\right)^{2k}\in O\big(\mathsf{poly}(m,n)\big)$ and $\Delta\eta\in O\bigg(\frac{1}{\sqrt{(1-\eta)^{n}}}\bigg)$. This indicates a regime of advantage for recycling mitigation over postselection up to a sample number
$N_{tot}\in O\bigg(\frac{\big(\frac{em}{n}\big)^{n}}{(1-\eta)^{n}\mathsf{poly}(m,n)}\bigg)\in e^{n\text{log}(\mathsf{poly}(n))+O(n)-\text{log}(\mathsf{poly(n))}+O(1)}.$ (5)
Consequently, recycling mitigation outperforms postselection for approximating the ideal probability up to an additive error
$\epsilon\in O\bigg(\mathsf{poly}(m,n)\Big(\frac{n}{m}\Big)^{n/2}\bigg)\in e^{-\frac{n}{2}\text{log}(\mathsf{poly}(n))+\text{log}(\mathsf{poly}(n))+O(1)}.$ (6)
Intuitively, because the statistical error for recycling mitigation is lower than that for postselection, the point up to which recycling mitigation outperforms postselection is typically determined by the bias error of the mitigation. Indeed, $O(\mathsf{poly}(m,n)\big(\frac{n}{m}\big)^{n/2})$ is a high-confidence upper bound on the bias error obtained in Theorem 10.
We present another recycling mitigation technique, exponential extrapolation, which outperforms the linear solving technique in numerical simulation experiments. Unlike linear solving, which uses a single value of $k$, exponential extrapolation uses recycled probabilities computed at multiple values of $k$, and attempts to extract ideal probabilities from the decay of these probabilities towards the uniform distribution (obtained at $k=n$ when all photons are lost). This decay is fitted to an exponential curve. The choice of the decay function is primarily heuristic and accurately captures the form of the decay function observed numerically.
We numerically observe that exponential extrapolation has lower bias error than linear solving. Although we do not show this analytically, we make the following conjecture about the performance of exponential extrapolation based on numerical evidence presented in Section 6.
Conjecture 3.
When $m\in\Omega(n^{2})$, and for values of $k\in O(1)$, exponential extrapolation recycling mitigation gives the following performance guarantee with high probability over the choice of $U$
$E_{s}(\epsilon_{{mitbias},s})\leq\delta(n)\Big(\frac{n}{m}\Big)^{n},$
where $E_{s}(\epsilon_{{mitbias},s})$ is the average, for fixed $U$, over all bit strings $s$ of the bias error $\epsilon_{{mitbias},s}$, and $\forall n$ $\delta(n)<1$ is some function of $n$ .
By using a Markov inequality, Conjecture 3 being true implies that, with high probability over the choice of $U$, and with probability $1-\frac{1}{w}$ over the choice of $s$ for a fixed $U$ and $w>1$, we have $\epsilon_{{mitbias},s}\leq w\delta(n)\big(\frac{n}{m}\big)^{n}$. In particular, one can choose a $w$ such that $w\delta(n)<1$, and relabel $\kappa(n)=w\delta(n)<1$, Conjecture 3 being true therefore implies that with high probability over $U$ and with probability $1-\frac{1}{w}$ over the choice of $s$ for a fixed $U$, $\epsilon_{{mitbias},s}\leq\kappa(n)\big(\frac{n}{m}\big)^{n}$ with $\kappa(n)<1$.
Furthermore, when $\eta$ is above some threshold value, we numerically find that the statistical error of extrapolation is lower than that of postselection. Let $\epsilon_{stat}(m,n,k,\eta,N_{tot})$ be the upper bound on the absolute value of the statistical error of extrapolation mitigation, we conjecture based on numerical evidence the following.
Conjecture 4.
For $\eta>\eta_{th}$, where $\eta_{th}\in[0,1]$ is some threshold loss value, we have that
$\epsilon_{stat}(m,n,k,\eta,N_{tot})\leq\frac{1}{\sqrt{(1-\eta)^{n}N_{tot}}}.$
Furthermore, for large enough $n$, $\epsilon_{stat}(m,n,k,\eta,N_{tot})\in o\bigg(\frac{1}{\sqrt{(1-\eta)^{n}N_{tot}}}\bigg)$.
If our conjectures on the bias and statistical errors of exponential extrapolation are true, we can show the following theorem in worst-case, when the errors on postselection and exponential extrapolation equal their upper bounds.
Theorem 5.
Assume that Conjectures 3 and 4 are true. Furthermore, assume that $n$ is sufficiently large and we are in the case where $|p_{mit}(s)-p_{id}(s)|=\epsilon_{stat}(m,n,k,\eta,N_{tot})+\kappa(n)\big(\frac{n}{m}\big)^{n}$, with $\kappa(n)<1$, and $|p_{post}(s)-p_{id}(s)|=\sqrt{\frac{1}{(1-\eta)^{n}N_{tot}}}$, where $p_{mit}(s)$ is the output probability of exponential extrapolation recycling mitigation. Then exponential extrapolation outperforms postselection for up to sample number $N_{tot}\ \in O\Big(\frac{(\frac{em}{n})^{2n}}{\kappa(n)^{2}(1-\eta)^{n}}\Big)\in e^{2n\log(\mathsf{poly}(n))+O(n)-\log(\kappa(n))+O(1)}$ and up to additive error $\epsilon\in O\Big(\kappa(n)\big(\frac{n}{m}\big)^{n}\Big)\in e^{-n\log(\mathsf{poly}(n))+\log(\kappa(n))+O(1)}$.
A proof of this is provided in Appendix 13.2.
Theorem 5 indicates that in worst-case, exponential extrapolation has better performance than linear solving, as the additive error up to which exponential extrapolation outperforms postselection is significantly smaller than that of linear solving.
In Section 5, we present the results of numerical simulations comparing the performance of the mitigation techniques with postselection. The recycling mitigation methods (linear solving and exponential extrapolation) consistently outperform postselection for a range of loss values and sample numbers. We also present results that indicates that in the average-case the errors of both recycling mitigation methods and postselection are below their analytically computed upper bounds. These experiments were performed in loss regimes relevant to near-term quantum photonics hardware.
In Section 6, we analyse the mitigated outputs in the regime $m\in\Omega(n^{2})$ for exponential extrapolation error mitigation, and conjecture that, in the absence of statistical errors, the distribution of $p_{mit}(s)$ is likely classically hard to compute. We support this conjecture with numerical calculations. Specifically, the conjecture is that computing the mitigated outputs is a problem that reduces to solving a restricted version of the $|GPE_{\pm}|^{2}$ problem [1]. This result is interesting because it indicates that, in general and in the absence of statistical error, the bias errors are small enough to render the error mitigated probabilities hard to compute classically.